What is the mass of 1.56 × 1021 atoms of magnesium in grams?
1
Expert's answer
2018-03-26T07:56:10-0400
Answer:
m (Mg) = n×M, where n is amount of substance, mole; M is molar mass, g/ mole
n (Mg) = N/Na, N is number of atoms, Na is Avogadro constant
1.56x10^21 atoms x (1 mol / 6.022x10^23 atoms) = 2.59x10^-3 moles

2.59x10^-3 moles x (24.305g / 1mol) = 0.630 grams Mg